### v2

# Corresponding author(s)

Abstract
DeepSeek-R1, known for its low training cost and exceptional reasoning capabilities, has achieved state-of-the-art performance on various benchmarks. However, detailed evaluations from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs. To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, and DeepSeek-R1-Distill-Llama series on the improved version A-Eval (A-Eval-2.0), an application-driven benchmark. By comparing original instruction-tuned models with their distilled counterparts, we analyze how reasoning enhancements impact performance across diverse practical tasks. Our results show that reasoning-enhanced models, while generally powerful, do not universally outperform across all tasks, with performance gains varying significantly across tasks and models. To further assist users in model selection, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models, ensuring optimal performance and resource efficiency in real-world applications. It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results. We will continuously optimize the evaluation benchmarks and periodically update this paper to provide more comprehensive and accurate evaluation results. Please refer to the latest version of the paper for the most recent results and conclusions.
Keywords:
DeepSeek application-driven evaluation reasoning enhancements capability boundary
1 Introduction
In recent years, large language models (LLMs) have revolutionized natural language processing (NLP). Models such as OpenAI’s GPT series [11], Alibaba’s Qwen series [14], MetaAI’s Llama series [3], and DeepSeek’s DeepSeek series [1] have not only advanced NLP technologies but also empowered intelligent solutions for real-world applications. Notably, DeepSeek-R1 [2], with its extremely low training cost, has achieved state-of-the-art (SOTA) performance comparable to OpenAI’s models [13]. Meanwhile, its exceptional reasoning capabilities have attracted significant attention.
Comprehensive evaluation is essential for understanding model strengths and limitations, as well as guiding optimization and practical deployment. While DeepSeek-R1 [2] has demonstrated SOTA performance on many existing benchmarks such as MMLU [5], C-Eval [6], SimpleQA [12], LiveCodeBench [7], Math-500 [10], these benchmarks do not fully capture the nuances of real-world applications. Consequently, two critical challenges arise when deploying DeepSeek models in practical scenarios: First, how should users determine the most suitable model scale? While larger models offer better capabilities, they also come with higher deployment and inference costs. Second, are reasoning-enhanced models always better for all tasks? If not, which tasks are best suited for reasoning-enhanced models, and which tasks can be handled by non-enhanced models?
To address the gap, we first create A-Eval-2.0 by modifying some QAs in A-Eval [9], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories. Then, we evaluate DeepSeek models on A-Eval-2.0. Through this evaluation, we extend existing evaluations and provide actionable insights into how reasoning enhancements influence model performance across various practical tasks. These results and findings help guide users in selecting the most suitable DeepSeek models for their specific needs, enabling cost-effective deployment in real-world applications.
It is important to emphasize that the selection of test samples and the design of evaluation criteria will inevitably introduce certain biases into the evaluation results. To address this, we will continue to optimize this evaluation work to enhance its comprehensiveness and reliability as much as possible.
2 Evaluation Framework
2.1 Evaluated Models
To assess the impact of reasoning enhancements in DeepSeek models for real-world applications, we evaluate 14 models grouped into seven pairs, including the Mixture-of-Experts (MoE) and dense models of various series and scales. Each pair consists of an instruction-tuned model and its corresponding distilled or reasoning-enhanced version, both derived from the same base model. Tab.  1 lists the evaluated models and their descriptions.
Table 1: The Evaluated model list.
Model Description
Qwen2.5-Math-1.5B-Instruct Dense math Model, SFT from Qwen2.5-Math-1.5B
DeepSeek-R1-Distill-Qwen-1.5B Dense model distilled from Qwen2.5-Math-1.5B
Qwen2.5-Math-7B-Instruct Dense math Model, SFT from Qwen2.5-Math-7B
DeepSeek-R1-Distill-Qwen-7B Dense model Distilled from Qwen2.5-Math-7B
Llama-3.1-8B-Instruct Dense model SFT from Llama-3.1-8B-Base
DeepSeek-R1-Distill-Llama-8B Dense model Distilled from Llama-3.1-8B-Base
Qwen2.5-14B-Instruct Dense model SFT from Qwen2.5-14B-Base
DeepSeek-R1-Distill-Qwen-14B Dense model Distilled from Qwen2.5-14B-Base
Qwen2.5-32B-Instruct Dense model SFT from Qwen2.5-32B-Base
DeepSeek-R1-Distill-Qwen-32B Dense model Distilled from Qwen2.5-32B-Base
Llama-3.3-70B-Instruct Dense model SFT from Llama-3.3-70B-Base
DeepSeek-R1-Distill-Llama-70B Dense model Distilled from Llama-3.3-70B-Instruct
DeepSeek-V3 (671B|37B) MoE model SFT from DeepSeek-V3-Base
DeepSeek-R1 (671B|37B) MoE model cold-start + RL from DeepSeek-V3-Base
2.2 Dataset
The A-Eval benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across five major categories and 27 subcategories. To accommodate the enhanced reasoning capabilities of the DeepSeek series models, we refine the QAs of Task Planning category in A-Eval, thereby constructing the evaluation dataset A-Eval-2.0 utilized in this evaluation. We use A-Eval-2.0’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications. This evaluation helps assess how DeepSeek-R1’s reasoning enhancements improve model capabilities in different practical scenarios.
2.3 Evaluation Process
We follow the automatic and manual scoring methods of A-Eval, and the automatic scoring process includes:
Inference: Feed each question ${{Q_{i}}}$ into the evaluated model to generate a prediction ${{P_{i}}}$.
Triplet Preparation: Construct the triplet $\left({{Q_{i}},{A_{i}},{P_{i}}}\right)$, where ${{A_{i}}}$ is the ground-truth answer for ${{Q_{i}}}$.
Scoring: Combine the prompt and triplet$\left({{Q_{i}},{A_{i}},{P_{i}}}\right)$, input them into the scoring model, and obtain the scoring output ${{S_{i}}}$ between 0 and 100. We use A-Eval’s scoring prompt and a greater LLM, Qwen2.5-72B-Instruct as the scoring model.
3 Results and Discussion
3.1 Overall Performance
Fig. 1 (a) presents the overall average scores of each model across all data.
Findings Consistent with Common Sense:
(1) Overall, reasoning-enhanced models outperform their original instruction-tuned counterparts.
(2) Overall, DeepSeek-V3 and DeepSeek-R1 demonstrate superior performance compared to other model families.
(3) Within the same series, models follow the scaling law [4, 8] regardless of whether they have undergone reasoning enhancement.
(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 169% and 53.6%, respectively. This is because they are math-focused and perform poorly on general tasks, but the distilled reasoning data from DeepSeek-R1 significantly enhances their general capabilities.
Figure 1: Average score of DeepSeek series models on A-Eval-2.0. (a) The overall average score of models across all data. (b) to (f). The average scores of models on each task. The "Instruct Model" refers to the original models without reasoning enhancement, and the "Reasoning Enhancement Model" refers to models that have been enhanced using DeepSeek-R1’s reasoning data.
Findings Contrary to Common Sense:
(5) Before and after distillation, Qwen2.5-32B outperforms the larger model, Llama-3.3-70B. This may be attributed to the fact that our evaluation dataset includes many Chinese data.
(6) Qwen-2.5-14B exhibits performance degradation after distillation.
3.2 Performance Comparison By Task
Fig. 1(b) to Fig. 1 (f) compare the scores of the models on the five major task categories.
Findings Consistent with Common Sense:
(7) The two mathematical models achieve comparable performance in Logical Reasoning and Task Planning but under-perform in other tasks.
(8) For Logical Reasoning tasks, all models show improvement after distillation, with the increase being greater than for other tasks.
Findings Contrary to Common Sense:
(9) Performance degradation occurs after distillation in: Text Understanding (Qwen2.5-14B, Qwen2.5-32B, and DeepSeek-V3), Information Extraction (Llama-3.1-8B, Qwen2.5-14B), and Text Generation (Qwen2.5-14B, DeepSeek-V3), and Task Planning (Llama-3.1-8B, Llama-3.3-70B).
(10) For Logical Reasoning tasks, original Llama-3.3-70B underperforms the smaller model Qwen2.5-14B. After distillation, DeepSeek-R1-Distilled-Llama-3.3-70B surpasses DeepSeek-R1-Distilled-Qwen2.5-32B.
(11) For Information Extraction tasks, Qwen2.5-32B consistently outperform the larger model Llama-3.3-70B both before and after distillation.
Figure 2: Average scores of the DeepSeek series models on the 27 subcategories.
3.3 Performance Comparison By Subtask
In more detail, Fig. 2 provides the scores of the models on the 27 subcategories.
After analyzing these results, here are some interesting findings:
(12) DeepSeek models dominate 23/27 subtasks, except for Short Text Classification, Named Entity Recognition, and Common Sense Question Answering.
(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-9.37), Part-of-Speech Tagging (-10.62), and Open-source Question Answering (-12.38%).
(14) Compared to other tasks, distillation brings the highest gains in Complex Mathematical Computation subtask, with an average improvement of 31.45%.
3.4 Performance Comparison By Model
To more clearly compare the performance of each group of models before and after reasoning enhancement on the five major tasks, we present the evaluation scores by seven model groups in Fig. 3.
Figure 3: Performance of each model group on five major tasks.
3.5 Model Selection Guidance for Users
To assist users in selecting the most suitable DeepSeek model based on their specific application requirements, we present the evaluation results using line charts, following the A-Eval methodology. These visualizations provide a clear and intuitive comparison of model performance across different tasks and subcategories.
Figure 4: Line chart for model selection on five major tasks.
Figure 4 illustrates the scores of the 14 evaluated models on each major task, as well as their average scores across all tasks. This comprehensive overview allows users to quickly identify models that excel in specific major task types or demonstrate strong overall performance.
In Fig. 5, we further break down the performance of the 14 models by showing their average scores within each of the five major task categories and their corresponding subcategories. These detailed visualizations enable users to assess model capabilities at a granular level, ensuring a precise match between model strengths and task requirements.
Figure 5: Line chart for model selection on five major tasks and corresponding subtasks.
Using these capability boundary quantization curves, users can make informed decisions when selecting the most suitable DeepSeek model for their specific tasks and performance expectations.
Table 2: Performance tier classification of DeepSeek models across major task categories. Models are categorized into tiers (A+: >85, A: 80-85, B: 75-80, C: 65-75, D: <65) based on their scores in Text Understanding (TU), Information Extraction (IE), Text Generation (TG), Logical Reasoning (LR), and Task Planning (TP).
Model Model Scale TU IE TG LR TP
Qwen2.5-Math-1.5B-Instruct 1.5B D D D D D
DeepSeek-R1-Distill-Qwen-1.5B 1.5B D D D D C
Qwen2.5-Math-7B-Instruct 7B D D D D D
DeepSeek-R1-Distill-Qwen-7B 7B C D D B C
Llama-3.1-8B-Instruct 8B C D C D B
DeepSeek-R1-Distill-Llama-8B 8B C D C C C
Qwen2.5-14B-Instruct 14B B B A B C
DeepSeek-R1-Distill-Qwen-14B 14B B C B A B
Qwen2.5-32B-Instruct 32B B B A B C
DeepSeek-R1-Distill-Qwen-32B 32B B B A B A
Llama-3.3-70B-Instruct 70B C C C C B
DeepSeek-R1-Distill-Llama-70B 70B B C B A C
DeepSeek-V3 671B|37B A B A+ A A+
DeepSeek-R1 671B|37B A A A A+ A+
To further simplify the model selection process for users without extensive knowledge of LLMs, we classify the performance of DeepSeek models into five tiers (A+, A, B, C, D) based on their scores in the five major task categories. As shown in Tab. 2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 65). This tiered classification provides an intuitive and accessible way for users to quickly identify the most suitable models for their specific application needs, balancing performance and cost-effectiveness.
Based on Tab. 2, we provide specific examples of model selection to illustrate its practical utility:
If users require A+ level performance, the following models are recommended for each task category:
Text Understanding: DeepSeek-V3 (A) or DeepSeek-R1 (A)
Information Extraction: DeepSeek-R1 (A)
Text Generation: DeepSeek-V3 (A+)
Logical Reasoning: DeepSeek-R1 (A+)
Task Planning: DeepSeek-V3 (A+) or DeepSeek-R1 (A+)
If users require A-level performance, the following models are recommended for each task category (highlighted in red in Tab. 2):
Text Understanding: DeepSeek-V3 (A)
Information Extraction: DeepSeek-R1 (A)
Text Generation: Qwen2.5-14B-Instruct (A)
Logical Reasoning: DeepSeek-R1-Distill-Qwen-14B (A)
Task Planning: DeepSeek-R1-Distill-Qwen-32B (A)
For users with more flexible requirements, such as B-level performance, the following models offer a cost-effective balance (highlighted in green in Tab. 2):
Text Understanding: Qwen2.5-14B-Instruct (B)
Information Extraction: Qwen2.5-14B-Instruct (B)
Text Generation: Qwen2.5-14B-Instruct (A)
Logical Reasoning: DeepSeek-R1-Distill-Qwen-7B (B)
Task Planning: DeepSeek-R1-Distill-Llama-8B (B)
These examples demonstrate how the tiered classification can guide users in selecting models that meet their specific performance and cost requirements, ensuring optimal deployment in real-world applications.
It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria. Therefore, our evaluation results may not be the most objective, comprehensive, or accurate. As we will continue to improve this evaluation work, we recommend referring to the latest version of the paper for the most recent experimental results and conclusions.
4 Conclusion
To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models and their distilled variants on the modified A-Eval-2.0 benchmark. Our analysis reveals that reasoning-enhanced models, while generally powerful, are not universally superior across all tasks. Finally, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models based on their specific application requirements. In the future, we will continue to advance this work by optimizing evaluation benchmarks and promptly updating evaluation results to the community.