### v3

# Corresponding author(s)

Abstract
DeepSeek-R1, known for its low training cost and exceptional reasoning capabilities, has achieved state-of-the-art performance on various benchmarks. However, detailed evaluations from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs. To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, DeepSeek-R1-Distill-Llama series, and their corresponding 4-bit quantized models on the enhanced A-Eval benchmark, A-Eval-2.0. By comparing original instruction-tuned models with their distilled counterparts, we analyze how reasoning enhancements impact performance across diverse practical tasks. Our results show that reasoning-enhanced models, while generally powerful, do not universally outperform across all tasks, with performance gains varying significantly across tasks and models. To further assist users in model selection, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models, ensuring optimal performance and resource efficiency in real-world applications. It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results. We will continuously optimize the evaluation benchmarks and periodically update this paper to provide more comprehensive and accurate evaluation results. Please refer to the latest version of the paper for the most recent results and conclusions.
Keywords:
DeepSeek application-driven evaluation reasoning enhancements capability boundary
1 Introduction
In recent years, large language models (LLMs) have revolutionized natural language processing (NLP). Models such as OpenAI’s GPT series [14], Alibaba’s Qwen series [18], MetaAI’s Llama series [4], and DeepSeek’s DeepSeek series [2] have not only advanced NLP technologies but also empowered intelligent solutions for real-world applications. Notably, DeepSeek-R1 [3], with its extremely low training cost, has achieved state-of-the-art (SOTA) performance comparable to OpenAI’s models [16]. Meanwhile, its exceptional reasoning capabilities have attracted significant attention.
Comprehensive evaluation is essential for understanding model strengths and limitations, as well as guiding optimization and practical deployment. While DeepSeek-R1 [3] has demonstrated SOTA performance on many existing benchmarks such as MMLU [6], C-Eval [7], SimpleQA [15], LiveCodeBench [8], Math-500 [13], these benchmarks do not fully capture the nuances of real-world applications. Consequently, three critical challenges arise when deploying DeepSeek models in practical scenarios: First, how should users determine the most suitable model scale? While larger models offer better capabilities, they also come with higher deployment and inference costs. Second, are reasoning-enhanced models always better for all tasks? If not, which tasks are best suited for reasoning-enhanced models, and which tasks can be handled by non-enhanced models? Third, is the extent of performance loss observed in the quantized model when applied to various real-world tasks within an acceptable range?
To address the gap, we first create A-Eval-2.0 by modifying some QAs in A-Eval [10], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories. Then, we evaluate DeepSeek models on A-Eval-2.0. Through this evaluation, we extend existing evaluations and provide actionable insights into how reasoning enhancements influence model performance across various practical tasks. These results and findings help guide users in selecting the most suitable DeepSeek models for their specific needs, enabling cost-effective deployment in real-world applications.
It is important to emphasize that the selection of test samples and the design of evaluation criteria will inevitably introduce certain biases into the evaluation results. To address this, we will continue to optimize this evaluation work to enhance its comprehensiveness and reliability as much as possible.
2 Evaluation Framework
2.1 Evaluated Models
To assess the impact of reasoning enhancements in DeepSeek models and assist users in identifying the most cost-effective models for practical applications, we conduct a systematic assessment of 21 models organized into seven distinct groups. These models encompass both Mixture-of-Experts (MoE) and dense models across various series and scales. Each group includes an instruction-tuned model, its corresponding distilled or reasoning-enhanced variant, and the associated quantized model. The instruction-tuned models are from HuggingFace  [17, 11], the distilled models are from HuggingFace  [1], and the quantized models are from Ollama  [12]. Table 1 lists the evaluated models and their descriptions.
Table 1: The Evaluated model list.
Model Description
Qwen2.5-Math-1.5B-Instruct Dense math Model, SFT from Qwen2.5-Math-1.5B
DeepSeek-R1-Distill-Qwen-1.5B Dense model distilled from Qwen2.5-Math-1.5B
Distill-Qwen-1.5B-Q4KM The 4-bit quantized version of distilled model
Qwen2.5-Math-7B-Instruct Dense math Model, SFT from Qwen2.5-Math-7B
DeepSeek-R1-Distill-Qwen-7B Dense model Distilled from Qwen2.5-Math-7B
Distill-Qwen-7B-Q4KM The 4-bit quantized version of distilled model
Llama-3.1-8B-Instruct Dense model SFT from Llama-3.1-8B-Base
DeepSeek-R1-Distill-Llama-8B Dense model Distilled from Llama-3.1-8B-Base
Distill-Llama-8B-Q4KM The 4-bit quantized version of distilled model
Qwen2.5-14B-Instruct Dense model SFT from Qwen2.5-14B-Base
DeepSeek-R1-Distill-Qwen-14B Dense model Distilled from Qwen2.5-14B-Base
Distill-Qwen-14B-Q4KM The 4-bit quantized version of distilled model
Qwen2.5-32B-Instruct Dense model SFT from Qwen2.5-32B-Base
DeepSeek-R1-Distill-Qwen-32B Dense model Distilled from Qwen2.5-32B-Base
Distill-Qwen-32B-Q4KM The 4-bit quantized version of distilled model
Llama-3.3-70B-Instruct Dense model SFT from Llama-3.3-70B-Base
DeepSeek-R1-Distill-Llama-70B Dense model Distilled from Llama-3.3-70B-Instruct
Distill-Llama-70B-Q4KM The 4-bit quantized version of distilled model
DeepSeek-V3 (671B|37B) MoE model SFT from DeepSeek-V3-Base
DeepSeek-R1 (671B|37B) MoE model cold-start + RL from DeepSeek-V3-Base
DeepSeek-R1-Q4KM The 4-bit quantized version of DeepSeek-R1
2.2 Dataset
The A-Eval [10] benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across three difficulty level, five major categories and 27 subcategories. To accommodate the enhanced reasoning capabilities of the DeepSeek series models, we refine the sub-tasks and QAs of Task Planning category in A-Eval, thereby constructing the evaluation dataset A-Eval-2.0 utilized in this evaluation. We use A-Eval-2.0’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications. On one hand, the evaluation on A-Eval-2.0 enables researchers to assess how the reasoning enhancements in DeepSeek-R1 improve model performance across various practical scenarios. On the other hand, the evaluation results assist users in selecting the most cost-effective models for their specific application needs.
2.3 Evaluation Process
We follow the automatic and manual scoring methods of A-Eval, and the automatic scoring process includes:
Inference: Feed each question ${{Q_{i}}}$ into the evaluated model to generate a prediction ${{P_{i}}}$.
Triplet Preparation: Construct the triplet $\left({{Q_{i}},{A_{i}},{P_{i}}}\right)$, where ${{A_{i}}}$ is the ground-truth answer for ${{Q_{i}}}$.
Scoring: Combine the prompt and triplet$\left({{Q_{i}},{A_{i}},{P_{i}}}\right)$, input them into the scoring model, and obtain the scoring output ${{S_{i}}}$ between 0 and 100. We use A-Eval’s scoring prompt and a greater LLM, Qwen2.5-72B-Instruct as the scoring model.
3 Results and Discussion
In this section, we first present the overall average performance of 21 models across all data in A-Eval-2.0 (Section 3.1). Next, we analyze the models’ performance from three perspectives: task category (Section 3.2), sub-task category (Section 3.3), and model group (Section 3.4). Based on these evaluation results, we provide line charts and tier classification table to assist users in selecting the most cost-effective models for their specific requirements (Section 3.5). Finally, we statistic performance of models on datasets with three difficulty levels—easy, medium, and hard—and discuss models’ performance to variations in question difficulty (Section 3.6).
3.1 Overall Performance
Figure 1 (a) presents the overall average scores of each model group across all data.
Findings Consistent with Common Sense:
(1) Undoubtedly, DeepSeek-V3 and DeepSeek-R1 demonstrate the best overall performance among all evaluated models..
(2) Overall, reasoning-enhanced models outperform their original instruction-tuned counterparts. After 4-bit quantization, the model’s scores decrease by an average of 3.52%.
(3) Within the same series, models follow the scaling law [5, 9] regardless of whether they have undergone reasoning enhancement.
(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 178.74% and 54.36%, respectively. This is because they are math-focused and perform poorly on general tasks, but the distilled reasoning data from DeepSeek-R1 significantly enhances their general capabilities.
Figure 1: Average score of DeepSeek series models on A-Eval-2.0. (a) The overall average score of models across all data. (b) to (f). The average scores of models on each task. The "Instruct Model" refers to the original Instruction-tuned models without reasoning enhancement, the "Reasoning Enhancement Model" refers to reasoning-enhanced models that have been enhanced using DeepSeek-R1’s reasoning data, and the "Quantization Model" refers to 4-bit quantized version of the reasoning enhancement models.
Findings Contrary to Common Sense:
(5) Both before and after distillation, Qwen2.5-32B consistently outperforms the larger Llama-3.3-70B model. The counterintuitive result is likely due to the predominance of Chinese data in our evaluation dataset, and this phenomenon also appears later in the discussion of performance comparison by task (Section 3.2) and by subtask (Section 3.3).
(6) Qwen-2.5-14B exhibits performance degradation after distillation.
3.2 Performance Comparison By Task
Figure 1(b) to  Figure 1 (f) compare the scores of the models on the five major task categories.
Findings Consistent with Common Sense:
(7) The two mathematical models achieve comparable performance in Logical Reasoning and Task Planning but under-perform in other tasks.
(8) For Logical Reasoning tasks, all models show improvement after distillation, with the increase being greater than for other tasks. Additionally, the performance degradation of 4-bit quantization on Logical Reasoning tasks is the most significant, at 7.91%.
Figure 2: Average scores of the DeepSeek series models on the 27 subcategories.
Findings Contrary to Common Sense:
(9) Performance degradation occurs after distillation in: Text Understanding (Qwen2.5-14B, Qwen2.5-32B, and DeepSeek-V3), Information Extraction (Llama-3.1-8B, Qwen2.5-14B), and Text Generation (Qwen2.5-14B, DeepSeek-V3), and Task Planning (Llama-3.1-8B, Llama-3.3-70B).
(10) For Logical Reasoning tasks, original Llama-3.3-70B underperforms the smaller model Qwen2.5-14B. After distillation, DeepSeek-R1-Distilled-Llama-3.3-70B surpasses DeepSeek-R1-Distilled-Qwen2.5-32B.
(11) For Information Extraction tasks, Qwen2.5-32B consistently outperform the larger model Llama-3.3-70B both before and after distillation.
3.3 Performance Comparison By Subtask
In more detail, Figure 2 provides the scores of the models on the 27 subcategories.
After analyzing these results, here are some interesting findings:
(12) DeepSeek models dominate 23/27 subtasks, except for Short Text Classification, Named Entity Recognition, and Common Sense Question Answering.
(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-11.3%), Part-of-Speech Tagging (-13.4%), and Open-source Question Answering (-17.05%).
(14) Compared to other tasks, distillation brings the highest gains in Complex Mathematical Computation subtask, with an average improvement of 31.45%.
3.4 Performance Comparison By Model
To more clearly compare the performance of each group of models before and after reasoning enhancement on the five major tasks, we present the evaluation scores by seven model groups in Figure 3.
Figure 3: Performance of each model group on five major tasks.
3.5 Model Selection Guidance for Users
To assist users in selecting the most suitable DeepSeek model based on their specific application requirements, we present the evaluation results using line charts, following the A-Eval methodology. These visualizations provide a clear and intuitive comparison of model performance across different tasks and subcategories.
Figure 4: Line chart for model selection on five major tasks.
Figure 4 illustrates the scores of the 21 evaluated models on each major task, as well as their average scores across all tasks. This comprehensive overview allows users to quickly identify models that excel in specific major task types or demonstrate strong overall performance.
In Figure 5, we further break down the performance of the 21 models by showing their average scores within each of the five major task categories and their corresponding subcategories. These detailed visualizations enable users to assess model capabilities at a granular level, ensuring a precise match between model strengths and task requirements.
Figure 5: Line chart for model selection on five major tasks and corresponding subtasks.
Using these capability boundary quantization curves, users can make informed decisions when selecting the most suitable DeepSeek model for their specific tasks and performance expectations.
Table 2: Performance tier classification of DeepSeek models across major task categories. Models are categorized into tiers (A+: >85, A: 80-85, B: 75-80, C: 65-75, D: <65) based on their scores in Text Understanding (TU), Information Extraction (IE), Text Generation (TG), Logical Reasoning (LR), and Task Planning (TP).
Model Model Scale TU IE TG LR TP
Qwen2.5-Math-1.5B-Instruct 1.5B D D D D D
DeepSeek-R1-Distill-Qwen-1.5B 1.5B D D D D C
Distill-Qwen-1.5B-Q4KM 1.5B D D D D C
Qwen2.5-Math-7B-Instruct 7B D D D D D
DeepSeek-R1-Distill-Qwen-7B 7B C D D B C
Distill-Qwen-7B-Q4KM 7B D D D C C
Llama-3.1-8B-Instruct 8B C D C D B
DeepSeek-R1-Distill-Llama-8B 8B C D C C C
Distill-Llama-8B-Q4KM 8B C D C D C
Qwen2.5-14B-Instruct 14B B B A B C
DeepSeek-R1-Distill-Qwen-14B 14B B C B A B
Distill-Qwen-14B-Q4KM 14B B C A C B
Qwen2.5-32B-Instruct 32B B B A B C
DeepSeek-R1-Distill-Qwen-32B 32B B B A B A
Distill-Qwen-32B-Q4KM 32B B B A+ B B
Llama-3.3-70B-Instruct 70B C C C C B
DeepSeek-R1-Distill-Llama-70B 70B B C B A C
Distill-Llama-70B-Q4KM 70B B C B C C
DeepSeek-V3 671B|37B A B A+ A A+
DeepSeek-R1 671B|37B A A A A+ A+
DeepSeek-R1-Q4KM 671B|37B B C A A+ A+
To further simplify the model selection process for users without extensive knowledge of LLMs, we classify the performance of DeepSeek models into five tiers (A+, A, B, C, D) based on their scores in the five major task categories. As shown in Table 2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 65). This tiered classification provides an intuitive and accessible way for users to quickly identify the most suitable models for their specific application needs, balancing performance and cost-effectiveness.
Based on Table 2, we provide specific examples of model selection to illustrate its practical utility:
If users require A+ level performance, the following models are recommended for each task category ((highlighted in blue in Table 2)):
Text Understanding: DeepSeek-V3 (A) or DeepSeek-R1 (A)
Information Extraction: DeepSeek-R1 (A)
Text Generation: Distill-Qwen-32B-Q4KM (A+)
Logical Reasoning: DeepSeek-R1-Q4KM (A+)
Task Planning: DeepSeek-R1-Q4KM (A+)
If users require A level performance, the following models are recommended for each task category (highlighted in red in Table 2):
Text Understanding: DeepSeek-V3 (A) or DeepSeek-R1 (A)
Information Extraction: DeepSeek-R1 (A)
Text Generation: Distill-Qwen-14B-Q4KM (A)
Logical Reasoning: DeepSeek-R1-Distill-Qwen-14B (A)
Task Planning: DeepSeek-R1-Distill-Qwen-32B (A)
For users with more flexible requirements, such as B level performance, the following models offer a cost-effective balance (highlighted in green in Table 2):
Text Understanding: Distill-Qwen-14B-Q4KM (B)
Information Extraction: Qwen2.5-14B-Instruct (B)
Text Generation: Distill-Qwen-14B-Q4KM (A)
Logical Reasoning: DeepSeek-R1-Distill-Qwen-7B (B)
Task Planning: Distill-Qwen-14B-Q4KM (B)
These examples demonstrate how the tiered classification can guide users in selecting models that meet their specific performance and cost requirements, ensuring optimal deployment in real-world applications.
3.6 Performance Comparison By Difficulty Level
In A-Eval and A-Eval-2.0, each QA pair is assigned a difficulty level label: Easy, Medium, or Hard. It is evident that models tend to achieve lower scores as the difficulty of the questions increases. However, there has been no quantitative analysis of how DeepSeek models perform on questions of varying difficulty levels. To address this gap, we separately analyze the evaluation results for each difficulty level and present our findings.
Figure 6 (a) illustrates the average scores of 21 models across task categories for Easy, Medium, and Hard data. Meanwhile, Figure 6 (b) - (f) depict the scores of the 21 models for each task category across the three difficulty levels.
Figure 6: Average performance of 21 models on Easy, Medium, and Hard data.
In Figure 7, we present the scores of 21 models on Easy, Medium, and Hard data for each sub-task category.
Figure 7: Performance of 21 models on Easy, Medium, and Hard data for each task and sub-task.
In Figure 8, we analyze the percentage score gains brought by reasoning enhancements on Easy, Medium, and Hard data. An interesting observation emerges:
(14) As the difficulty of the data increases, distillation or reasoning enhancements yield higher performance gains for the models. Interestingly, reasoning enhancements can even have a negative impact on simpler problems.
Figure 8: Performance gain of reasoning enhancement on Easy, Medium, and Hard data. (a) Performance gain of each model group across all tasks. (b) Performance gain of each task category across all models (excluding math models).
It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria. Therefore, our evaluation results may not be the most objective, comprehensive, or accurate. As we will continue to improve this evaluation work, we recommend referring to the latest version of the paper for the most recent experimental results and conclusions.
4 Conclusion
To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models, their distilled variants, and 4-bit quantized models on the modified A-Eval-2.0 benchmark. Our analysis reveals that reasoning-enhanced models, while generally powerful, are not universally superior across all tasks. Finally, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models based on their specific application requirements. In the future, we will continue to advance this work by optimizing evaluation benchmarks and promptly updating evaluation results to the community.